# Evaluating Meta-Regression Techniques: A Simulation Study on Heterogeneity in Location and Time

### v1

@@ 5.1.1 Bivariate Data Generation Process - 1 Covariate
$\beta_{1}=0.5=(1)^{-1}(0.5)$
$\sigma^{2}=0.75=1-(0.5)(1)^{-1}(0.5)$
@@ 5.1.2 Multivariate Data Generation Process - 3 or 5 Covariates
For the multivariate covariate cases (three or five), the variance-covariance matrix changes compared to Section 5.1.1, so the true parameter values also change. Equation 20 shows the updated joint distribution for three covariates, Equation 21 shows the updated joint distribution for five covariates, and both equations show how this changes the values of the true parameters. The details of the data-generating process, details of the joint distribution, and the reading of the elements of the joint distribution below are detailed in Section 5.1.1 above.
@@ 5.2 Performance Criteria
In the literature on meta-analysis simulations, there is variation on which criteria are considered to be evidence of optimal model performance, typically due to the particular measure or method that is the focus of the research. This study includes the following as performance criteria that dictate which of the tested methods performed optimally in the following hierarchy: an optimal balance of power (i.e., neither under-powered nor over-powered), estimator bias and the precision of confidence intervals. Typically, confidence intervals are only included when new confidence interval estimators are being developed, but we include this since it can capture precision of the estimate (Pateras et al., 2018; Langan et al., 2019; Van Aert & Jackson, 2019; Noma et al., 2020; Richardson et al., 2021). Precision of variance in estimates are not included in the performance criteria of this manuscript since this is not a typical performance criteria in the literature. This is because variance is already captured as an element of accuracy of the estimate via mean squared error (MSE) as $MSE()=Var()+Bias()^{2}$ (Blázquez-Rincón et al., 2023). Note that in the meta-analysis simulation literature, it is rare for any goodness of fit measures to be used as performance criteria. The only exception found in our literature review was one seminal study that used Akaike Information Criteria (AIC) (T. Stanley et al., 2023). We accordingly carefully considered whether it might be appropriate for our simulation or not, but do not include any goodness of fit measures for the following reasons. First, $R^{2}$ is designed differently for random and fixed effects, and the way $R^{2}$ is measured is different as well. When testing our base cases, we observed stark differences in $R^{2}$, even when all of our other performance criteria were identical between random and fixed effects). To add further complications, $\bar{R}^{2}$ would need to be used instead since the number of covariates is one of our parameters that vary. Second, the simulation literature does not typically consider information criteria since it does not explicitly capture model fit, it only captures whether more variables should be added to the model or not by adding a penalty to each variable added (Akaike, 1974; Sugasawa & Noma, 2021; Pateras et al., 2018). Its consistency has also been disproven in simulations (Spanos, 2010).
@@
Our performance criteria is made up of three measures assessed in a particular order. Power is assessed first because if an estimator has insufficient power, it is very likely estimator bias will increase and predictive accuracy will decrease (Gambarota & Altoè, 2024; Pateras et al., 2018; Schwarzer et al., 2024; Spanos, 2019a; Wooldridge, 2010). When reviewing our results we saw exactly this; when power was poor, all of the other performance criteria uniformly performed substantially worse. This is because if an estimator has insufficient power to detect discrepancies from the null hypothesis, then it is more likely to be biased. This bias harms the accuracy of coefficient estimates, which accordingly harms the confidence intervals, model robustness, and predictive accuracy. Suppose, however, that power is strong for all methodology specifications, then estimator bias would be assessed next to determine which specification provides the most accurate effect estimation. If all of the above performance criterion are equal, then confidence intervals are compared to assess the precision of the estimates (Pateras et al., 2018; Langan et al., 2019; Van Aert & Jackson, 2019; Noma et al., 2020; Richardson et al., 2021). The remaining subsections will review each of the performance criteria used in more detail.
@@ 5.2.1 Statistical Power
The most dominant performance criterion for this simulation is an optimal balance of statistical power (i.e. neither underpowered nor overpowered). Power to detect discrepancies from the null hypothesis (also known as 1-Type II error) is a common performance criterion in the meta-analysis simulation literature, it is typically measured as the percent of iterations with statistically significant results in the simulation literature (Gambarota & Altoè, 2024; Pateras et al., 2018; Schwarzer et al., 2024). To measure power more comprehensively, this study calculated power by finding the area under the non-central t distribution to the left of the Type I error (Cameron & Trivedi, 2005). In addition, the figures in this manuscript show the power to detect a range of discrepancies from the null hypothesis via the x-axis (Lehmann & Romano, 2005). This goes further than typical power measures since it ensures the effects of small and large discrepancies from the null hypothesis on the estimator’s power can be precisely captured (i.e. small and large effect sizes respectively) (Spanos, 2019b; D’Orazio, 2023; Greene, 2018; Stock, 2020). The estimators power increases when the discrepancy from the null hypothesis increases and the power decreases when the discrepancy decreases. The Type I error was adjusted when the sample size of the simulated meta-regression was increased by either a larger sample size or more countries. This was done to combat the Large N problem which has been shown to overpower models when the sample size is too large, which can lead to erroneous inferences due to erroneous results that appear statistically significant (Spanos, 2023; Myors & Murphy, 2023). Testing the power of each methodology specification with a variety of parameters shows the capacity of each methodology tested in the simulation to detect discrepancies from the null, and how each parameter can affect this capacity. For instance, if one of the ten methodology specifications has much lower power than the others when there is a large time effect ceterus paribus, then it can be inferred that one methodology does not have as much power when large time-level heterogeneity is present.
@@ 5.2.2 Estimator Bias
If power is equal between methodologies, the next performance criteria to consider is the estimator bias of the covariates. Estimator bias is captured by which methodology specification provides the smallest MSE value (mean squared error) through $\frac{1}{n}\sum^{N}_{i=1}(\beta_{i}-\hat{\beta}_{i})^{2}$. This measure is commonly included in the simulation literature because it measures how close the estimates are to actual values as seen in Section 5.1 (Bergtold et al., 2010; Spanos & McGuirk, 2001; Dormann et al., 2020; Blázquez-Rincón et al., 2023; Poom & Af Wåhlberg, 2022; Erez et al., 1996). The combinations of parameters and the varying levels of location and time heterogeneity specified in Section 5.1 allow us to see how precise the estimates of the coefficients are for each methodology tested in the simulation. For instance, if one of the ten specifications has a much higher MSE than the others when there is a large location effect ceteris paribus, then it can be inferred that one methodology does not handle large magnitudes of location-level heterogeneity well when estimating the covariate.
There are other variations of MSE which are rare in the meta-analysis simulation literature, but well circulated in statistics and econometrics literature including mean absolute error, mean percent error, or mean absolute percent error (respectively MAE, MPE or MAPE) (Gneiting, 2011; West, 2006; Timmermann, 2006; Maronna et al., 2019; Hyndman & Athanasopoulos, 2021). While MSE is designed to penalize larger deviations from the true value, MAE is designed without penalty via $\frac{1}{n}\sum^{N}_{i=1}|\beta_{i}-\hat{\beta}_{i}|$. Section 5.1 shows that our simulation is not designed to have outliers, but if it did, MAE would be more appropriate to use than MSE due to the sensitivity of MSE to outliers (Maronna et al., 2019; Müller & Guido, 2017; Hyndman & Athanasopoulos, 2021; French et al., 1987). MSE is scale dependent, so if this is a concern it may be more appropriate to use MPE which is scale independent and measures the percentage difference via $\frac{1}{n}\sum^{N}_{i=1}\frac{\beta_{i}-\hat{\beta}_{i}}{\beta_{i}}$ (Gneiting, 2011; West, 2006; Diebold & Lopez, 1996). MAPE is simply the absolute value of the percentage difference from MPE which in addition to being scale independent it focuses on the magnitude of prediction error instead of over/under prediction like MPE (Gneiting, 2011; West, 2006; Timmermann, 2006). The simulation of this manuscript will primarily use MSE, but the other variations may be considered if MSE values are equal for all cases.
5.2.3 Residual Diagnostic
If covariate estimator bias are equal between methodology specifications, the next performance criteria to consider is the residual diagnostic via the residuals MSE through $\frac{1}{n}\sum^{N}_{i=1}(Y_{i}-\hat{Y}_{i})^{2}$. MSE of the residual is commonly included in the simulation literature because this measures how close $\hat{Y}$ is to $Y$, which reveals how precise the estimate predictions of the variable of interest are to the actual data (Bergtold et al., 2010; Richardson et al., 2021; McGrath et al., 2019; Langan et al., 2019; Blázquez-Rincón et al., 2023; Sidik & Jonkman, 2007). Note that even though we have full information by generating the simulation data with a pre-specified variance-covariance matrix, diagnosing the residual is still important to see if anything is not able to be captured by the model. For instance, if one of the ten specification has a much higher MSE than the others when there is a large time effect ceteris paribus, then it can be inferred that one method does not handle large magnitudes of time-level heterogeneity well. This section can be improved by considering more formal tests for residual diagnostics beyond testing the means difference from zero. We plan to approach this by testing if the IID assumption holds (i.e. independence and constant variance). As mentioned in Section 5.2.2 the simulation of this manuscript will primarily use MSE, but the other variations may be considered if MSE values are equal for all cases. The equation for MAPE is the same, and the equations for MAE and MPE are slightly different as shown below
$MAE=\frac{1}{n}\sum^{N}_{i=1}|Y_{i}-\hat{Y}_{i}|,\quad\quad MPE=\frac{1}{n}\sum^{N}_{i=1}\frac{Y_{i}-\hat{Y}_{i}}{Y_{i}}$
@@ 5.2.4 Confidence Interval
If power and estimator bias are equal between methodology specifications, the next performance criteria to consider are confidence intervals. This has been extensively covered in the meta-analysis simulation literature with the focus of capturing estimate precision (Pateras et al., 2018; Langan et al., 2019; Van Aert & Jackson, 2019; Noma et al., 2020; Richardson et al., 2021). The particular focus is primarily to ensure the coefficient estimate is within the estimated confidence interval, and then the focus is on the confidence interval having the smallest spread possible (i.e. being as precise as possible). For instance, even if all methodology specifications include the true value in their coefficient estimate, a method that has the smallest range in its confidence interval would be considered the best (in terms of which specification has the most precise estimate).
@@ 5.2.5 Effects of Parameters on Performance Criteria
This subsection will review the common effects that variations in the simulation parameters have on each of the performance criteria according to the meta-analysis simulation literature. The magnitude of power increases when sample size is larger with all other parameters equal (Blázquez-Rincón et al., 2023; Richardson et al., 2021; Kulinskaya et al., 2021; Noma et al., 2020; Spanos, 2019a; Wooldridge, 2010). This increase in power due to sample size increasing also decreases estimator bias (Gambarota & Altoè, 2024; Pateras et al., 2018; Schwarzer et al., 2024; Spanos, 2019a; Wooldridge, 2010). In comparison, the magnitude of power decreases when sample size is small with all other parameters equal which increase estimator bias. In the simulation design, the number of studies and locations are linked to the sample size since the number of studies is driven by the number of locations, and when there are more locations the sample size of the meta-regression is larger. The magnitude of power increases with the number of covariates being higher with all other parameters equal (Stock, 2020; Greene, 2018; Lehmann & Romano, 2005). This increase in power from more covariates decreases the estimator bias (Gambarota & Altoè, 2024; Pateras et al., 2018; Schwarzer et al., 2024; Spanos, 2019a; Wooldridge, 2010). In contrast, when the number of covariates is small, the magnitude of power decreases and estimator bias increases. To recap, the effects of parameters on our performance criteria have been well documented in the meta-analysis simulation literature, these can be seen in Table 3 below.
Power Estimator Bias
$\uparrow n$uparrow $\downarrow$
$\downarrow n$downarrow $\uparrow$
$\uparrow$ Covariates $\uparrow$ $\downarrow$
$\downarrow$ Covariates $\downarrow$ $\uparrow$
@@ 5.3 Results
It can be seen in the twelve cases for the ten methodology specifications that each specification performs slightly better when it has more covariates (excluding $RE_{l}$, which does not gain power). Figures 4 - 6 in Appendix C below show performance criteria from a sample result to illustrate this one versus five covariate comparison (ceteris paribus, see the supplementary material for all the results). Figures 4 & 16 in Appendix C reveal power nor the estimator bias differ for any of the methodology specifications, but Figure 5 reveals that the residuals MSE is uniformly lower for all specifications. By conducting residual diagnostics with MSE it can be seen if any methodology specifications have issues with capturing anything systematic in the data (i.e. model robustness). As mentioned in Section 5.2.3, we plan to further improve these residual diagnostics by testing if the IID assumption holds (i.e. independence and constant variance). Note that the two specifications that model time heterogeneity with a trend term add value by quantifying the overall trend over time (rather than being simply controlled). Figure 6 in Appendix C shows no difference between the 1 and 5 covariate case. Note that the estimates for the covariate coefficients are accurate for all methodology specifications for each case, highlighting why solely coefficient accuracy is an insufficient performance criterion.
@@
To summarize our overall simulation results beyond the number of covariates, the standard methodology performed the best and the specifications that control joint heterogeneity performed the second best. This includes the $FE_{l,Trend}$ specification we introduced in this manuscript, but excludes $FE_{s,Trend}$ proposed in the meta-analysis simulation literature since it provided inaccurate trend estimates. The worst-performing specifications were those that did not jointly control for location and time heterogeneity (including location-level fixed effects proposed in the experimental economics literature). Further details are broken down into the following subsections: Section 5.3.1 covers the methodology specifications that are best controlled for time heterogeneity, as well as the impact of a small time effect (compared to a large time effect). Section 5.3.2 covers the impact of a small location effect (compared to a large location effect), while Section 5.3.3 covers the impact when the number of locations differ. Lastly, Section 5.3.4 briefly summarizes the results and highlights the key takeaways. For further details, please see the supplementary materials which include both a table with the mean values from all iterations and histograms that show all values from all iterations.
5.3.1 Impact of Small versus Large Time Effect

### v2

@@ 5.1.1 Bivariate Data Generation Process - 1 Covariate
$\beta_{1}=0.5=(1)^{-1}(0.5),\quad\quad\sigma^{2}=0.75=1-(0.5)(1)^{-1}(0.5)$
@@ 5.1.2 Multivariate Data Generation Process - 3 or 5 Covariates
For the multivariate covariate cases (three or five), the variance-covariance matrix changes compared to Section 5.1.1, so the true parameter values also change. Equation 11 shows the updated joint distribution for three covariates, Equation 12 shows the updated joint distribution for five covariates, and both equations show how this changes the values of the true parameters. Both equations can be found in Appendix C. The details of the data-generating process, details of the joint distribution, and the reading of the elements of the joint distribution below are detailed in Section 5.1.1 above.
@@ 5.2 Performance Criteria
In the literature on meta-analysis simulations, there is variation on which criteria are considered to be evidence of optimal model performance, typically due to the particular measure or method that is the focus of the research. This study includes the following as performance criteria that dictate which of the tested methods performed optimally in the following hierarchy: an optimal balance of power (i.e., neither under-powered nor over-powered), estimator bias, the precision of variance, and the precision of confidence intervals. Typically, confidence intervals are only included when new confidence interval estimators are being developed, but we include this since it can capture precision of the estimate (Pateras et al., 2018; Langan et al., 2019; Van Aert & Jackson, 2019; Noma et al., 2020; Richardson et al., 2021). Although precision of variance is not typically included in the performance criteria of meta-analysis simulation literature, we include it in our performance criteria since mean square error (MSE) is regularly used to either measure bias, or complement another bias measure (Richardson et al., 2021; McGrath et al., 2019; Langan et al., 2019; T. D. Stanley & Doucouliagos, 2017). Variance and bias are both elements of MSE via $MSE()=Var()+Bias()^{2}$, but we disentangle them from MSE to assess both the estimator bias as well as the variance in the estimates (Blázquez-Rincón et al., 2023). Note that if the estimator is unbiased, then $MSE\approx Var()$ since $Bias\approx 0$. Goodness of fit measures are rarely used in the meta-analysis simulation literature, the only exception found in our literature review was one seminal study that used Akaike Information Criteria (AIC) (T. Stanley et al., 2023). We accordingly carefully considered whether it might be appropriate for our simulation or not, but do not include any goodness of fit measures for the following reasons. First, $R^{2}$ is designed differently for random and fixed effects, and the way $R^{2}$ is measured is different as well. When testing our base cases, we observed stark differences in $R^{2}$, even when all of our other performance criteria were identical between random and fixed effects). To add further complications, $\bar{R}^{2}$ would need to be used instead since the number of covariates is one of our parameters that vary. Second, the simulation literature does not typically consider information criteria since it does not explicitly capture model fit, it only captures whether more variables should be added to the model or not by adding a penalty to each variable added (Sugasawa & Noma, 2021; Pateras et al., 2018; Akaike, 1974). Its consistency has also been disproven in simulations (Spanos, 2010).
@@
Our performance criteria is made up of four measures assessed in a particular order. Power is assessed first because if an estimator has insufficient power, it is very likely estimator bias will increase and predictive accuracy will decrease (Gambarota & Altoè, 2024; Schwarzer et al., 2024; Spanos, 2019a; Pateras et al., 2018; Wooldridge, 2010). When reviewing our results we saw exactly this; when power was poor, all of the other performance criteria uniformly performed substantially worse. This is because if an estimator has insufficient power to detect discrepancies from the null hypothesis, then it is more likely to be biased. This bias harms the accuracy of coefficient estimates, which accordingly harms the confidence intervals, model robustness, and predictive accuracy. Suppose, however, that power is strong for all methodology specifications, then estimator bias would be assessed next to determine which specification provides the most accurate effect estimation. This is commonly used in the meta-regression simulation literature, but there are variations on how this is measured, so we define the measure we use in Section 5.2.2. If both power and biases of the estimator are equal, then estimator variance would be assessed next. Lastly, if all of the above performance criterion are equal, then confidence intervals are compared to assess the precision of the estimates (Richardson et al., 2021; Van Aert & Jackson, 2019; Noma et al., 2020; Pateras et al., 2018; Langan et al., 2019).
@@ 5.2.1 Statistical Power
The most dominant performance criterion for this simulation is an optimal balance of statistical power (i.e. neither underpowered nor overpowered). Power to detect discrepancies from the null hypothesis (also known as 1-Type II error) is a common performance criterion in the meta-analysis simulation literature, it is typically measured as the percent of iterations with statistically significant results in the simulation literature (Gambarota & Altoè, 2024; Schwarzer et al., 2024; Pateras et al., 2018). To measure power more comprehensively, this study calculated power by finding the area under the non-central t distribution to the left of the Type I error (Cameron & Trivedi, 2005). In addition, the figures in this manuscript show the power to detect a range of discrepancies from the null hypothesis via the x-axis (Lehmann & Romano, 2005). This goes further than typical power measures since it ensures the effects of small and large discrepancies from the null hypothesis on the estimator’s power can be precisely captured (i.e. small and large effect sizes respectively) (D’Orazio, 2023; Stock, 2020; Spanos, 2019b; Greene, 2018). To calculate the power to measure varying discrepancies from the null hypothesis, a series of alternate hypotheses $\beta_{H_{A}}^{(i)}=\beta+0.1\cdot i$ for $i=(1,2,...,100)$ are used. Then, the non centrality parameter $\delta^{(i)}$ is calculated for each alternate hypothesis, which is then used to calculate the power of the estimator for each alternate hypothesis as follows in the below equation:
$Power^{(i)}=1-\mathbb{P}(t_{df,\delta^{(i)}}<c)\quad s.t.\quad\delta^{(i)}=\frac{|\beta-\beta_{H_{A}}^{(i)}|}{\hat{\sigma}}$ (10)
The test statistic is represented as $t$, the critical value is denoted as $c$, and $\hat{\sigma}$ represents the estimators standard error. The estimators power increases when the discrepancy from the null hypothesis increases and the power decreases when the discrepancy decreases. The Type I error was adjusted when the sample size of the simulated meta-regression was increased by either a larger sample size or more countries. This was done to combat the Large N problem which has been shown to overpower models when the sample size is too large, which can lead to erroneous inferences due to erroneous results that appear statistically significant (Spanos, 2023; Myors & Murphy, 2023). Testing the power of each methodology specification with a variety of parameters shows the capacity of each methodology tested in the simulation to detect discrepancies from the null, and how each parameter can affect this capacity. For instance, if one of the ten methodology specifications has much lower power than the others when there is a large time effect ceterus paribus, then it can be inferred that one methodology does not have as much power when large time-level heterogeneity is present.
@@ 5.2.2 Estimator Bias
If power is equal between methodologies, the next performance criteria to consider is the estimator bias of the covariates. Estimator bias is captured by which methodology specification provides the smallest discrepancy from the true values. This measure is the most dominant performance criteria in the simulation literature because it measures how close the estimates are to actual values, but is measured in various ways in the literature. We accordingly use the following measure $Bias=\frac{1}{N}\sum_{i=1}^{N}\hat{\beta}_{i}-\beta$ from the literature where $N$ is the total number of iterations in the Monte Carlo simulation, but we additionally use $Bias_{i}=\hat{\beta_{i}}-\beta$ to capture variation in each iteration within the simulation (Blázquez-Rincón et al., 2023; López‐López et al., 2014). The combinations of parameters and the varying levels of location and time heterogeneity specified in Section 5.1 allow us to see how accurate the estimates of the coefficients are for each methodology tested in the simulation. For instance, if one of the ten specifications has a much higher bias than the others when there is a large location effect ceteris paribus, then it can be inferred that one methodology does not handle large magnitudes of location-level heterogeneity well when estimating the covariate. Note that this is not the same thing as mean squared error (MSE) not only because the penalty is squared, but also because it captures both bias and variance in one combined measure (see Appendix H for further details on MSE and similar measures).
5.2.3 Estimator Variance
When both power and estimator bias are equal between methodologies, the next performance criteria to consider is the variance of the estimator. This is calculated as $Var=\sigma^{2}=(\sqrt{n}\cdot\hat{\sigma})^{2}$ from the equation $\hat{\sigma}=\frac{\sigma}{\sqrt{n}}$ (Van Aert & Wicherts, 2023; Wooldridge, 2023; Loucks & Van Beek, 2017; López‐López et al., 2014; Higgins et al., 2009). An ideal estimator will have minimal variance, which means that there will be minimal variation in the values of the estimate. This is important to capture since an estimator with minimal variance is indicative of stability and consistency of estimated values (Wooldridge, 2010; Cameron & Trivedi, 2005; Baltagi, 2005). In fact, estimates can be highly distorted with large variance even if their estimator is unbiased. This is especially important for data sets with smaller sample sizes. Note that variance is independent of accuracy of the estimate, it is possible for an estimate to not be accurate (i.e. biased) but have a small variance or for an estimate to be accurate (i.e. unbiased) but have a large variance. Also note that this is not necessarily identical to MSE because MSE captures both bias and variance in one combined measure.
To illustrate how this performance criteria will be used in assessing our simulation results, suppose one of the ten specifications has a much larger variance than the others when there is a large time effect ceteris paribus. It can then be inferred that one methodology does not handle large magnitudes of time-level heterogeneity well since the variation in the estimate is higher. This inference is valid even if the estimators bias is not different, which highlights why measuring bias and variance separately provides more information than MSE. In the meta-analysis simulation literature, there are measures for between-study variance, but since most of the methodology specifications use fixed effects rather than random effects, we do not use them. Instead, the standard variance measure shown above is employed since group level heterogeneity is controlled for in each methodology specification.
@@ 5.2.4 Confidence Interval
If power, estimator bias and variance are equal between methodology specifications, the next performance criteria to consider are confidence intervals. This has been extensively covered in the meta-analysis simulation literature with the focus of capturing estimate precision (Richardson et al., 2021; Van Aert & Jackson, 2019; Noma et al., 2020; Pateras et al., 2018; Langan et al., 2019). The primary focus is to ensure the coefficient estimate is within the estimated confidence interval, and then the focus is on the confidence interval having the smallest spread possible (i.e. highest precision). For instance, even if all methodology specifications include the true value in their coefficient estimate, a method that has the smallest range in its confidence interval would be considered best (since it is the most precise estimate).
@@ 5.2.5 Effects of Parameters on Performance Criteria
This subsection will review the common effects that variations in the simulation parameters have on each of the performance criteria according to the meta-analysis simulation literature. The magnitude of power increases when sample size is larger with all other parameters equal (Blázquez-Rincón et al., 2023; Richardson et al., 2021; Kulinskaya et al., 2021; Noma et al., 2020; Spanos, 2019a; Wooldridge, 2010). This increase in power due to sample size increasing also decreases estimator bias and variance (Gambarota & Altoè, 2024; Schwarzer et al., 2024; Spanos, 2019a; Pateras et al., 2018; Wooldridge, 2010). In comparison, the magnitude of power decreases when sample size is small with all other parameters equal which increase estimator bias and variance. In the simulation design, the number of studies and locations are linked to the sample size since the number of studies is driven by the number of locations, and when there are more locations the sample size of the meta-regression is larger. The magnitude of power increases with the number of covariates being higher with all other parameters equal (Stock, 2020; Greene, 2018; Lehmann & Romano, 2005). This increase in power from more covariates decreases the estimator bias and variance (Gambarota & Altoè, 2024; Schwarzer et al., 2024; Spanos, 2019a; Pateras et al., 2018; Wooldridge, 2010). In contrast, when the number of covariates is small, the magnitude of power decreases and estimator bias increases and variance . To recap, the effects of parameters on our performance criteria have been well documented in the meta-analysis simulation literature, these can be seen in Table 3 below.
@@ 5.3 Results
Figures 4 - 7 in Appendix D show performance criteria from a sample result to illustrate what happens when more covariates are added by comparing one versus five covariates (ceteris paribus). It can be seen for all cases and methodology specifications that each specification performs slightly better when it has more covariates. Figure 4 reveals that power does slightly increase, and Figures 6 & 7 reveal that although the estimator bias does not change, the estimator variance does uniformly lower for all specifications with more covariates. Note that the two specifications that model time heterogeneity with a trend term add value by quantifying the overall trend over time (rather than being simply controlled). Figure 5 in Appendix D shows no difference between the one and five covariate case. Note that the estimates for the covariate coefficients are accurate for all methodology specifications for each case, highlighting why solely coefficient accuracy is an insufficient performance criterion.
@@
To summarize our overall simulation results beyond the number of covariates, the standard methodology performed the best and the specifications that control joint heterogeneity performed the second best. This includes the $FE_{l,Trend}$ specification we introduced in this manuscript, but excludes $FE_{s,Trend}$ proposed in the meta-analysis simulation literature since it consistently provided biased trend estimates. The worst-performing specifications were those that did not jointly control for location and time heterogeneity (including location-level fixed effects proposed in the experimental economics literature). Note that for all cases the mean estimator was found to be unbiased, but at the Monte Carlo iteration level there were small levels of positive of negative bias. Accordingly, this means that on average $MSE\approx Var$, and increases/decreases in estimator bias discussed in this section are relative. Further details are broken down into the following subsections: Section 5.3.1 covers the methodology specifications that are best controlled for time heterogeneity, as well as the impact of a small time effect (compared to a large time effect). Section 5.3.2 covers the impact of a small location effect (compared to a large location effect), while Section 5.3.3 covers the impact when the number of locations differ. Lastly, Section 5.3.4 briefly summarizes the results and highlights the key takeaways.
5.3.1 Impact of Smaller versus Larger Time Effect